# An accurate mean-field equation for voter model dynamics on scale-free networks

### v1

@@ Abstract
Understanding the emergent macroscopic behavior of dynamical systems on networks is a crucial but challenging task. One of the simplest and most effective methods to construct a reduced macroscopic model is given by mean-field theory. The resulting approximations perform well on dense and homogeneous networks but poorly on scale-free networks, which, however, are more realistic in many applications. In this paper, we introduce a modified version of the mean-field approximation for voter model dynamics on scale-free networks. This approach achieves a significantly smaller approximation error than standard methods without increasing dimensionality.
@@ I Introduction
The voter model [1, 2] and its many extensions (see, e.g., [3, 4, 5] for overviews) are among the most popular models for studying spreading or infection processes on networks. In this paper, we focus on a general variant, the continous-time noisy voter model (CNVM) [6, 7, 8]. Given a simple graph $G$ on $N$ nodes, each node $i\in\{1,\dots,N\}$ has a state $x_{i}\in\{0,\dots,M-1\}=:[M]$. As the voter model is often applied to opinion dynamics, we refer to $x_{i}$ as the opinion of node $i$. Each opinion $x_{i}$ evolves over time according to a continuous-time Markov chain: if $x_{i}$ is in state $a\in[M]$, it transitions to a different state $b\in[M]$ at the rate
@@
An important observable for such discrete-state systems is the vector of shares $c\in[0,1]^{M}$, $\lVert c\rVert_{1}=1$, where $c_{a}$ describes the fraction of nodes $i$ that have opinion $x_{i}=a$. Using mean-field theory [9], it has been shown that, in the large-population limit $N\to\infty$, the trajectory $c(t)$ of time-dependent shares converges to the solution of a mean-field equation (MFE), provided that the underlying network is sufficiently homogeneous and dense [6, 10, 11, 12]. For scale-free networks—e.g., generated by the Barabási–Albert (BA) model [13] with preferential attachment parameter mm—however, this classical mean-field approximation performs poorly [14, 15, 16]. The main contribution of this paper is the derivation and validation of a modified mean-field equation (mMFE) that accurately captures the macroscopic behavior of the CNVM on BA networks.
@@
It is well known [6] that for the CNVM the MFE is a simple system of ordinary differential equations
$\frac{\text{d}}{\text{d}t}c(t)=\sum_{\begin{subarray}{c}a,b\in[M]\\ a\neq b\end{subarray}}c_{a}(t)\big{(}r_{a,b}\,c_{b}(t)+\tilde{r}_{a,b}\big{)}(e_{b}-e_{a}),$ (2)
@@
where $e_{a}$ denotes the $a$-th unit vector in $\mathbb{R}^{M}$. To achieve an accurate approximation for BA networks, we introduce two modifications. First, the imitation rates $r_{a,b}$ are multiplied by a constant correlation factor $0<\alpha\leq 1$; second, the shares $c\in[0,1]^{M}$ are replaced with the degree-weighted shares $\bar{c}\in[0,1]^{M}$, defined as the average over the states weighted by node degrees. The dynamics of these weighted shares is then approximated by the mMFE
$\frac{\text{d}}{\text{d}t}\bar{c}(t)=\sum_{a\neq b}\bar{c}_{a}(t)\big{(}\alpha\,r_{a,b}\,\bar{c}_{b}(t)+\tilde{r}_{a,b}\big{)}(e_{b}-e_{a}).$ (3)
@@
The significance of the degree-weighted shares is already known in some contexts [17, 18], and they were specifically determined to be excellent collective variables for the CNVM on BA networks using a data-driven method developed by the authors [19].
@@
In this paper, we derive the above mMFE (3), motivating why these modifications are necessary for BA networks, and show how to obtain the correlation factor $\alpha$. We analyze the quality and robustness of the mMFE depending on the model parameters $r$ and $\tilde{r}$, using a theoretical error analysis as well as numerical studies. We show that the mMFE approximates the degree-weighted shares exceptionally well when the imitation rates $r_{a,b}$ are similarly large and there is moderate noise $\tilde{r}$. Moreover, in this setting the mMFE is robust in the sense that $\alpha$ does not depend on the specific model parameters and initial state. In other parameter regimes, the approximation quality degrades slightly. We show that the mMFE outperforms standard methods like MFE and pair-approximation (PA) [4, 15] significantly. Another interesting result is that the correlation factor $\alpha$ increases towards $1$ with increasing preferential attachment parameter $m$ of the BA network.
@@
Accompanying code for reproducing the numerical experiments presented in this paper is provided at [20].
@@ II The modified mean-field equation
In this section, we derive the modified mean-field equation (mMFE) from the evolution equation of the first-order moments. We have to rely on two assumptions, the correlation assumption and the interchangeability assumption, which are explained in more detail later. For simplicity, we consider the case of $M=2$ opinions. The general case is analogous, and is treated in the appendix A. In the binary-state case, it is sufficient to only work with the (expected) weighted share $\bar{c}_{1}$ of opinion $1$, which we will denote as $\bar{c}=\bar{c}_{1}$ for simplicity. It can be defined as
$\bar{c}:=\mathbb{E}\Big{[}\frac{1}{\sum_{i}d_{i}}\sum_{i\in[N]}d_{i}x_{i}\Big{]}=\frac{1}{\sum_{i}d_{i}}\sum_{i\in[N]}d_{i}\mathbb{E}[x_{i}],$ (4)
where expectation is taken with respect to stochasticity in the dynamical evolution and possibly uncertainty in the initial state. Recall that in the CNVM the rate at which node $i$ in state $x_{i}=0$ transitions to state $1$ is given by
@@
where $A$ denotes the adjacency matrix of the network. Hence, the evolution of the first-order moment $\mathbb{E}[x_{i}]$ is
@@
$\mathbb{P}(x_{i}=0)\ \mathbb{E}\Big{[}\frac{r_{0,1}}{d_{i}}\sum_{j}A_{i,j}x_{j}+\tilde{r}_{0,1}\mid x_{i}=0\Big{]}\$-
$\mathbb{P}(x_{i}=1)\ \mathbb{E}\Big{[}\frac{r_{1,0}}{d_{i}}\sum_{j}A_{i,j}(1-x_{j})+\tilde{r}_{1,0}\mid x_{i}=1\Big{]}.$ (6)
@@
for some constant correlation factor $0<\alpha<1$. This implies that, on average, it is more likely to find two neighbors in the same state than if they were stochastically independent. Given this assumption, we have
@@
which is a weaker form of stating that all nodes have the same average state $\bar{c}$, and are therefore interchangeable. Inserting yields the mMFE for $M=2$ opinions
@@
It is clear that, due to possible violation of the two assumptions, the solution of the mMFE (3) does not exactly match the actual weighted shares $\bar{c}$. To distinguish the two, we will continue to refer to the true weighted shares as $\bar{c}$ and denote the solution of the mMFE as $\bar{c}_{\text{{\tiny mMFE}}}$.
@@ II.1 Relation to classical mean-field theory
Moreover, in the classical theory it is typically argued that nodes become interchangeable in the mean-field limit, so that the expected state of every node is simply given by the average state, e.g., $\mathbb{E}[x_{i}]=c$ for all $i$ in the binary-state case ($M=2$). We observe numerically that, for scale-free networks, this interchangeability is also valid on average, although the expected state of each node is related to the degree-weighted average state $\bar{c}$ instead, as described in assumption (12). As mentioned earlier, this finding was inspired by the authors’ work [19], which revealed computationally that the weighted shares are excellent collective variables for the CNVM on BA networks.
@@ II.2 Error analysis
$L_{1}(T;\alpha):=\frac{1}{T}\int_{0}^{T}\frac{1}{2}\lVert\bar{c}(t)-\bar{c}_{\text{{\tiny mMFE}}}(t)\rVert_{1}\text{d}t.$ (19)
(Note that $\frac{1}{2}\lVert\bar{c}(t)-\bar{c}_{\text{{\tiny mMFE}}}(t)\rVert_{1}\in[0,1]$ is the total variation distance between discrete probability distributions [21].) Then, choosing the factor
@@
yields the optimal fit of $\bar{c}_{\text{{\tiny mMFE}}}$ to $\bar{c}$ on the interval $[0,T]$.
@@ II.3 Parameter regimes
    Regime D: Dissimilar rates and (partially) no noise. Example: $r_{0,1}=1$, $r_{1,0}=0$, $\tilde{r}_{0,1}=0$, $\tilde{r}_{1,0}=0.3$ (i.e., SIS model [22]).
(The case of large noise is not considered, as it would render the impact of the underlying network negligible.)
We outline the differences between these regimes using the above-mentioned example parameters (rescaled to exhibit comparable behavior over the interval $t\in[0,100]$). In all regimes, we use the same initial state, in which a randomly chosen $10\%$ of nodes start with opinion 1 and the remainder with opinion 0, and the same BA network with parameter $m=3$ and $N=10^{4}$ nodes. Fig. 1 shows, for each regime, the errors $\delta_{\text{cor}}$ and $\delta_{\text{int}}$ for the optimal $\alpha^{*}$ as defined in (20) with $T=100$, estimated from 15000 model simulations, as well as the resulting $L_{1}$-qerror (19). Fig. 2 (left) shows the corresponding trajectories $\bar{c}(t)$ and $\bar{c}_{\text{{\tiny mMFE}}}(t)$.
Figure 2: Trajectories of $\bar{c}$, estimated from simulations, for the example systems of the different regimes A B,C,D (left), compared to the mMFE (II). The mMFE uses the optimal $\alpha^{*}$ as in (20), shown on the right (green dashed). The right plot also shows the $\alpha$ that minimizes $\delta_{\text{cor}}(t;\alpha)$ for each $t$ (orange solid).
@@
Figure 3: Interchangeability error $\delta_{\text{int}}$ when using the regular shares $c$ instead of the degree-weighted shares $\bar{c}$ in (15), for regimes A and B (with different initial conditions than in Fig. 1).
It should also be noted that the use of the degree-weighted shares $\bar{c}$, as opposed to the regular shares $c$, is crucial for obtaining a small error in the initial phase. Fig. 3 illustrates that using the regular shares $c$ in the interchangeability assumption (12) instead of $\bar{c}$ can result in an error $\delta_{\text{int}}$ that is magnitudes larger. This is demonstrated using an initial state $x$ where the top $10\%$ of highest-degree nodes have opinion $1$ and the rest opinion 0, resulting in a large discrepancy between $c$ and $\bar{c}$.
@@ II.4 Practical considerations
Computing the factor $\alpha^{*}$ by solving (20) can be computationally demanding as many simulations may be required to estimate the weighted shares trajectory $\bar{c}(t)$, $t\in[0,T]$. However, $\alpha^{*}$ can be calculated efficiently using only the steady-state weighted shares $\bar{c}_{\infty}$, which can be obtained from a single (sufficiently long) simulation started from an arbitrary initial state. Consider that in the steady-state of the mMFE (3) it is
@@
for all $a\in[M]$, where summation is performed over all $b$ except $b=a$. This implies that
@@
Hence, $\alpha$ is given by inserting $\bar{c}_{\infty}$ into (22). In the ergodic setting, e.g., if $\tilde{r}_{a,b}>0$ for all $a\neq b$, the above equation yields the optimal $\alpha^{*}$ as defined in (20) with $T\to\infty$.
The usefulness of the mMFE would be severely limited if $\alpha$ had to be calculated from (20) or (22) for every system of interest. However, the next section shows that this is not necessary as the optimal $\alpha$ is robust across initial states and a wide range of system parameters.

### v3

@@ Abstract
Understanding the emergent macroscopic behavior of dynamical systems on networks is a crucial but challenging task. One of the simplest and most effective methods to construct a reduced macroscopic model is given by mean-field theory. The resulting approximations perform well on dense and homogeneous networks but poorly on scale-free networks, which, however, are more realistic in many applications. In this paper, we introduce a modified version of the mean-field approximation for voter model dynamics on scale-free networks. The two main deviations from classical theory are that we use degree-weighted shares as coarse variables and that we introduce a correlation factor that can be interpreted as slowing down dynamics induced by interactions. We observe that the correlation factor is only a property of the network and not of the state or of parameters of the process. This approach achieves a significantly smaller approximation error than standard methods without increasing dimensionality.
@@ I Introduction
The voter model [Clifford1973, Holley1975] and its many extensions (see, e.g., [Castellano2009, Porter2016, Redner2019] for overviews) are among the most popular models for studying spreading or infection processes on networks. In this paper, we focus on a general variant, the continous-time noisy voter model (CNVM) [Luecke2023, Luecke2024a, Carro2016]. Given a simple graph $G$ on $N$ nodes, each node $i\in\{0,\dots,N-1\}=:[N]$ has a state $x_{i}\in\{0,\dots,M-1\}=:[M]$. As the voter model is often applied to opinion dynamics, we refer to $x_{i}$ as the opinion of node $i$. Each opinion $x_{i}$ evolves over time according to a continuous-time Markov chain: if $x_{i}$ is in state $a\in[M]$, it transitions to a different state $b\in[M]$ at the rate
@@
An important observable for such discrete-state systems is the vector of shares $\bm{c}=(c_{a})_{a\in[M]}\in[0,1]^{M}$, $\lVert\bm{c}\rVert_{1}=1$, where $c_{a}$ describes the expected fraction of nodes $i$ that have opinion $x_{i}=a$, i.e.,
$c_{a}:=\frac{1}{N}\sum_{i\in[N]}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}],$ (2)
@@
where $\mathbf{1}_{\{x_{i}=a\}}\in\{0,1\}$ denotes the indicator function of the event $x_{i}=a$. Using mean-field theory [Barrat2008], it has been shown that, in the large-population limit $N\to\infty$, the trajectory $\bm{c}(t)$ of time-dependent shares converges to the solution of a mean-field equation (MFE), provided that the underlying network is sufficiently homogeneous and dense [Luecke2023, Keliger2024, Gleeson2012, Ayi2021]. For scale-free networks—e.g., generated by the Barabási–Albert (BA) model [Albert2002] with preferential attachment parameter mm—however, this classical mean-field approximation performs poorly [Fernley2022, Gleeson2013, PastorSatorras2001]. The main contribution of this paper is the derivation and validation of a modified mean-field equation (mMFE) that can accurately capture the macroscopic behavior of the CNVM on scale-free networks.
@@
It is well known [Luecke2023] that for the CNVM the MFE is a simple system of ordinary differential equations
$\frac{\text{d}}{\text{d}t}\bm{c}(t)=\sum_{\begin{subarray}{c}a,b\in[M]\\ a\neq b\end{subarray}}c_{a}(t)\big(r_{a,b}\,c_{b}(t)+\tilde{r}_{a,b}\big)(\bm{e}_{b}-\bm{e}_{a}),$ (3)
@@
where $\bm{e}_{a}$ denotes the $a$-th unit vector in $\mathbb{R}^{M}$. To achieve an accurate approximation for scale-free networks, we introduce two modifications. First, the imitation terms $r_{a,b}c_{a}c_{b}$ are multiplied by a constant correlation factor $0<\alpha\leq 1$; second, the shares $\bm{c}\in[0,1]^{M}$ are replaced by the degree-weighted shares $\bar{\bm{c}}=(\bar{c}_{a})_{a\in[M]}\in[0,1]^{M}$,
$\bar{c}_{a}:=\frac{1}{\sum_{i\in[N]}d_{i}}\sum_{i\in[N]}d_{i}\mathbb{E}[\mathbf{1}_{\{x_{i}=a\}}].$ (4)
The dynamics of these degree-weighted shares is then approximated by the mMFE
$\frac{\text{d}}{\text{d}t}\bar{\bm{c}}(t)=\sum_{\begin{subarray}{c}a,b\in[M]\\ a\neq b\end{subarray}}\bar{c}_{a}(t)\big(\alpha\,r_{a,b}\,\bar{c}_{b}(t)+\tilde{r}_{a,b}\big)(\bm{e}_{b}-\bm{e}_{a}).$ (5)
@@
The significance of the degree-weighted shares is already known in some contexts [Bianconi2002, Suchecki2005], and they were specifically determined to be excellent collective variables for the CNVM on BA networks using a data-driven method developed by the authors [Luecke2024].
@@
In this paper, we derive the above mMFE (5), motivating why these modifications are necessary for scale-free networks, and show how to obtain the correlation factor $\alpha$. We analyze the quality and robustness of the mMFE depending on the model parameters $r$ and $\tilde{r}$, using a theoretical error analysis as well as numerical studies. We show that the mMFE approximates the degree-weighted shares exceptionally well when the imitation rates $r_{a,b}$ are similarly large and there is moderate noise $\tilde{r}$. Moreover, in this setting the mMFE is robust in the sense that $\alpha$ does not depend on the specific model parameters and initial state—hence, the correlation factor appears to be a universal attribute of the network, more precisely of the network’s statistical properties. In other parameter regimes, the approximation quality degrades slightly. Due to its popularity, the numerical examples presented in the main text utilize the BA model to generate scale-free networks, but almost identical results can be obtained for other examples like the configuration model [Bollobas1980], as shown in Appendix B. Specifically for BA networks, another interesting result is that the correlation factor $\alpha$ increases towards $1$ with increasing preferential attachment parameter $m$ of the BA network. Finally, we show that the mMFE outperforms standard methods of similar complexity like the standard homogeneous MFE (3) and pair-approximation (PA) [Porter2016, Gleeson2013] significantly. While higher-dimensional methods like the heterogeneous MFE and PA [Porter2016, Pugliese2009], approximate master equation [gleeson2011high], quenched MFE and PA [Mata2013], or dynamic cavity method [Neri2009] are able to resolve more detailed information about the system than the mMFE, they also require substantially more variables. The mMFE is, on the other hand, able to provide a parsimonious macroscopic description of the collective dynamics using only a minimal set of variables: the degree-weighted shares $\bar{\bm{c}}$.
@@
Accompanying code for reproducing the numerical experiments presented in this paper is provided at [lucke_2025_17122500].
@@ II The modified mean-field equation
In this section, we derive the modified mean-field equation (mMFE) from the evolution equation of the first-order moments. We have to rely on two assumptions, the correlation assumption and the interchangeability assumption, which are explained in more detail later. For simplicity, we consider the case of $M=2$ opinions. The general case is analogous, and is treated in Appendix A. In the binary-state case, it is sufficient to only work with the (expected) weighted share $\bar{c}_{1}$ of opinion $1$, which we will denote as $\bar{c}=\bar{c}_{1}\in[0,1]$ for simplicity. It can be defined as
$\bar{c}:=\mathbb{E}\bigg[\frac{1}{\sum_{i}d_{i}}\sum_{i\in[N]}d_{i}x_{i}\bigg]=\frac{1}{\sum_{i}d_{i}}\sum_{i\in[N]}d_{i}\mathbb{E}[x_{i}],$ (6)
where expectation is taken with respect to randomness in the dynamical evolution and possibly also in the initial state. Recall that in the CNVM the rate at which node $i$ in state $x_{i}=0$ transitions to state $1$ is given by
@@
where $(A_{i,j})_{i,j\in[N]}\in\{0,1\}^{N\times N}$ denotes the adjacency matrix of the network. Hence, the evolution of the first-order moment $\mathbb{E}[x_{i}]$ is
@@
$\mathbb{P}(x_{i}=0)\ \mathbb{E}\Big[\frac{r_{0,1}}{d_{i}}\sum_{j}A_{i,j}x_{j}+\tilde{r}_{0,1}\,\Big|\,x_{i}=0\Big]\$-
$\mathbb{P}(x_{i}=1)\ \mathbb{E}\Big[\frac{r_{1,0}}{d_{i}}\sum_{j}A_{i,j}(1-x_{j})+\tilde{r}_{1,0}\,\Big|\,x_{i}=1\Big].$ (8)
@@
for some constant correlation factor $0<\alpha<1$. This implies that, on average, it is more likely to find two neighbors in the same state than if they were stochastically independent (in which case $\alpha=1$)  [^1] This is because if $x_{i}$ and $x_{j}$ were stochastically independent, we would have that $\mathbb{E}[(1-x_{i})x_{j}]$ $=\mathbb{P}[x_{i}=0,\ x_{j}=1]$ $=\mathbb{P}[x_{i}=0]\mathbb{P}[x_{j}=1]=\mathbb{E}[1-x_{i}]\mathbb{E}[x_{j}].$ Hence, the weak correlation assumption states that, on network-average, it is less probable than in the independent case that neighbors have different opinions.. Given this assumption, we have
@@
which states that the average pairwise interaction between the expected states behaves as if all nodes shared a common state $\bar{c}$. This is substantially weaker than a “strong” interchangeability assumption of the form $\mathbb{E}[x_{i}]=\zeta$ for all nodes $i$, which would imply $c=\bar{c}=\zeta$. Applying (14) yields the mMFE for $M=2$ opinions
@@
It is clear that, due to possible violation of the two assumptions, the solution of the mMFE (5) does not exactly match the actual weighted shares $\bar{\bm{c}}$. To distinguish the two, we will continue to refer to the true weighted shares as $\bar{\bm{c}}$ and denote the solution of the mMFE as $\bar{\bm{c}}_{\text{{\tiny mMFE}}}$.
@@ II.1 Relation to classical mean-field theory
Moreover, in the classical theory it is typically argued that nodes become interchangeable in the mean-field limit, so that the expected state of every node is simply given by the average state, e.g., $\mathbb{E}[x_{i}]=c$ for all $i$ in the binary-state case ($M=2$). We observe numerically that, for BA networks, this interchangeability is also valid on average; however, although the expected state of each node is related to the degree-weighted average state $\bar{c}$ instead, as described in assumption (14). As mentioned earlier, this finding was inspired by the authors’ work [Luecke2024], which computationally revealed that the weighted shares are excellent collective variables for the CNVM on BA networks. This means that their future evolution depends only on the current value of $\bar{c}$, and on no further information from the full state $\bm{x}=(x_{i})_{i\in[N]}$.
@@ II.2 Error analysis
$L_{1}(T;\alpha):=\frac{1}{T}\int_{0}^{T}\frac{1}{2}\lVert\bar{\bm{c}}(t)-\bar{\bm{c}}_{\text{{\tiny mMFE}}}(t)\rVert_{1}\text{d}t.$ (21)
(Note that $\frac{1}{2}\lVert\bar{\bm{c}}(t)-\bar{\bm{c}}_{\text{{\tiny mMFE}}}(t)\rVert_{1}\in[0,1]$ is the total variation distance between discrete probability distributions [Rachev1991].) Then, choosing the factor
@@
yields the optimal fit of $\bar{\bm{c}}_{\text{{\tiny mMFE}}}$ to $\bar{\bm{c}}$ on the interval $[0,T]$.
@@ II.3 Parameter regimes
    Regime D: Dissimilar rates and (partially) no noise. Example: $r_{0,1}=1$, $r_{1,0}=0$, $\tilde{r}_{0,1}=0$, $\tilde{r}_{1,0}=0.3$ (i.e., SIS model [Kiss2017]).
(The case of large noise is not considered, as it would render the impact of the underlying network negligible.) A possible intuitive interpretation is that regimes A and B describe opinion or strategy dynamics with bidirectional imitation and some exploration—regime A being nearly neutral, regime B allowing moderate bias. Regime C reflects pure imitation without exploration (e.g., consensus formation), while regime D captures strongly asymmetric, directional contagion (e.g., disease or innovation spreading). Note that the particular values $0.05$ and $0.5$ that serve as the similarity thresholds for regimes A and B were chosen as useful rules of thumb but do not carry a special meaning. In fact, the mMFE error grows steadily with the rate dissimilarity, see Fig. S3 in the appendix for details.
We outline the differences between these regimes using the above-mentioned example parameters (rescaled to exhibit comparable behavior over the interval $t\in[0,100]$). In all regimes, we use the same initial state, in which a randomly chosen $10\%$ of nodes start with opinion 1 and the remainder with opinion 0, and the same BA network with parameter $m=3$ and $N=10^{4}$ nodes. Fig. 1 shows, for each regime, the errors $\delta_{\text{cor}}$ and $\delta_{\text{int}}$ for the optimal $\alpha^{*}$ as defined in (22) with $T=100$, estimated from 15000 model simulations, as well as the resulting $L_{1}$-error (21). Fig. 2 (left) shows the corresponding trajectories $\bar{c}(t)$ and $\bar{c}_{\text{{\tiny mMFE}}}(t)$.
Figure 2: Trajectories of $\bar{c}$, given in (6), estimated from simulations, for the example systems of the different regimes A B,C,D (left), compared to the mMFE (II). The mMFE uses the optimal $\alpha^{*}$ as in (22), shown on the right (green dashed). The right plot also shows the $\alpha$ that minimizes $\delta_{\text{cor}}(t;\alpha)$ for each $t$ (orange solid).
@@
Figure 3: Interchangeability error $\delta_{\text{int}}$, see (17), when using the regular shares $c$ instead of the degree-weighted shares $\bar{c}$, for regimes A and B (with different initial conditions than in Fig. 1).
It should also be noted that the use of the degree-weighted shares $\bar{c}$, as opposed to the regular shares $c$, is crucial for obtaining a small error in the initial phase. Fig. 3 illustrates that using the regular shares $c$ in the interchangeability assumption (14) instead of $\bar{c}$ can result in an error $\delta_{\text{int}}$ that is magnitudes larger. This is demonstrated using an initial state $\bm{x}$ where the top $10\%$ of the highest-degree nodes have opinion $1$ and the rest have opinion 0, resulting in a large discrepancy between $c$ and $\bar{c}$. This behavior reflects the disproportionate dynamical influence of high-degree nodes in the CNVM: hubs are imitated more often and affect many neighbors simultaneously, which strongly shapes the early and transient dynamics. The degree-weighted shares $\bar{c}$ encode this effect and therefore satisfy the closure assumptions much better than the unweighted shares $c$, as Fig. 3 illustrates.
@@ II.4 Practical considerations
Computing the factor $\alpha^{*}$ by solving (22) can be computationally demanding, as many simulations may be required to estimate the weighted shares trajectory $\bar{\bm{c}}(t)$, $t\in[0,T]$. In the ergodic system case, i.e., if $\tilde{r}_{a,b}>0$ for all $a\neq b$, the factor $\alpha^{*}$ can be calculated efficiently using only the steady-state weighted shares $\bar{\bm{c}}_{\infty}=(\bar{c}_{\infty,a})_{a\in[M]}$, which can be obtained from a single (sufficiently long) simulation started from an arbitrary initial state. Consider that in the steady-state of the mMFE (5) it is
@@
for all $a\in[M]$, where the summation is performed over all $b$ except $b=a$. This implies that
@@
Hence, $\alpha$ is given by inserting $\bar{\bm{c}}_{\infty}$ into (24). Note that, due to the ergodicity, it is $0<\bar{c}_{\infty,a}<1$ for all $a$, and hence the above expression is well-defined and yields the optimal $\alpha^{*}$ as defined in (22) with $T\to\infty$.
The usefulness of the mMFE would be severely limited if $\alpha$ had to be calculated from (22) or (24) for every system of interest. However, the next section shows that this is not necessary, as the optimal $\alpha$ is robust across initial states and a wide range of system parameters.